€250 is invested in a savings account. The nominal rate value of interest convertiblemonthly for the first three months is 18%,and the nominal rate of interest convertible quarterly for the next 9 months is 20%.
How much is in the account at the end of the year
1
Expert's answer
2020-04-21T08:18:55-0400

S=P*(1+j/m)^N

The nominal rate is an annual rate that accrues interest m times a year (m > 1). Denote it by j . Therefore, for one period, interest is accrued at the j / m rate.

N is the total number of accrual periods.

j/m=1.5% in three months (18/12),

j/m=5% in quarterly (20%/4),

9 months is three quarters.

Sum of Deposit=P*(1+0.015)³*(1+0.05)³,

S=P*1.21,

S=250*1.21=302.5 .